# Prospects of constraining on the polarizations of gravitational waves from binary black holes using space- and ground-based detectors

### v1

@@ Abstract
In general relativity, gravitational waves (GWs) exhibit two tensor modes, while alternative theories predict up to six polarization modes. We investigate GW polarization constraints using a model-independent parametrized post-Einsteinian framework with space- and ground-based detectors. Evaluating LISA, Taiji, TianQin, LIGO, Virgo, KAGRA, and Einstein Telescope (ET), we analyze their capabilities and network performance. Taiji provides the best constraints among space-based detectors, with LISA outperforming TianQin. Among ground-based detectors, LIGO excels in vector modes, while ET offers the most comprehensive constraints. In network scenarios, LISA+TJm performs best, and ET surpasses second-generation detector combinations. Multiband observations alleviate scalar mode degeneracies, significantly enhancing ground-based detector performance. Combined space- and ground-based observations yield robust constraints on GW polarization, advancing tests of deviations from general relativity. Our findings highlight the potential of future GW missions to refine our understanding of gravitational physics through precise polarization measurements.
@@ I Introduction
Over the past century, general relativity (GR) has been rigorously tested through numerous solar system experiments and astrophysical observations, with no definitive evidence suggesting deviations from GR [1, 2, 3]. Despite passing all experimental tests, there are indications that GR may need to be extended, particularly due to the challenges posed by dark energy and dark matter in cosmology [4, 5]. Gravitational waves (GWs) offer a new avenue for testing GR, especially in the strong-field regime where GWs are generated by compact objects [6]. One method to test GR is by detecting additional polarizations of GWs. In GR, GWs only have two tensor modes, whereas some alternative gravity theories predict up to six polarization modes [7]. For instance, Brans-Dicke theory predicts an additional scalar mode [2, 8], while $f(R)$ theory and Horndeski theory propose two additional scalar modes [9, 10, 11]. Einstein-Aether theory predicts five polarization modes [12, 13], and some tensor-vector-scalar theories encompass all six polarization modes [14, 15]. Therefore, detecting additional polarizations of GWs can help identify deviations from GR, surpass current experimental limits, and reveal the deeper nature of gravity.
@@
Since LIGO first detected GW [16], nearly a hundred GW events generated by compact binary coalescences (CBCs) have been observed [17, 18, 19]. Testing GR with these GW data has shown that all observed data are consistent with GR so far [20, 21, 22, 23, 24, 25, 26]. With the commencement of the fourth observation run (O4), more GW events are expected to be detected, allowing for further testing of GR. In general, to detect the polarization of transient GWs, more detectors are needed to obtain responses from different directions, thereby enhancing detection capabilities [27]. Ground-based detectors in the hertz frequency band, such as LIGO [28], Virgo [29], and KAGRA [30], collaborate to form a detector network to detect the polarization of GWs. For continuous GWs emitted by distant CBCs, a detector with directional change capability is sufficient [27]. Space-based GW detectors in the millihertz frequency band, such as LISA [31], Taiji [32], and TianQin [33], consist of three spacecraft orbiting the Sun/Earth in a triangular formation, making it possible to detect additional GW polarizations due to the detector’s motion in space [34].
@@
Besides determining whether the detector can theoretically detect additional polarizations of GWs, it is also crucial to ascertain whether the detected GW signal is consistent with the predictions of GR. One approach is to use a waveform model with purely phenomenological parameters to represent possible deviations from GR and to apply constraints to these parameters using observed GW data [6]. Yunes and Pretorius proposed the parameterized post-Einstein (ppE) waveform [35], based on the post-Newtonian (PN) approximation, which is suitable for parameterizing the influence of alternative gravity theories on non-GR polarizations. Most modified gravity theories, such as Brans-Dicke theory, massive gravity, and bimetric theory, can be described using the ppE framework [36]. Although the ppE framework cannot parameterize all possible deviations from GR, it can test for additional polarizations, providing a method to evaluate the performance of detectors in testing GR [37, 38, 39].
@@
Currently, some work utilizes the ppE framework to constrain additional polarizations for testing GR. Narikawa and Tagoshi discussed the potential of advanced ground-based GW detectors, such as LIGO, Virgo, and KAGRA, in detecting general deviations between GWs and the predictions of GR using the ppE framework [40]. Huwyler et al. employed LISA to detect massive black hole binaries (MBHBs) and test GR by representing ppE waveforms with phase corrections only [41]. References [36, 6] discussed the potential of Taiji and TianQin in using MBHBs to test GR and provide constraints on non-GR parameters. For detector networks, Nair et al. focused on the synergistic effect of the Einstein Telescope (ET) and preDECIGO, demonstrating enhanced sensitivity within a specific band [42]. In Ref. [43], Wang and Han used the LISA-Taiji network to test the ppE parameters of deviations from the GR waveform, indicating a significant improvement in the detection of polarization amplitude through joint observations.
@@
Based on our previous work [44, 45, 46], we investigate the constraints from binary black holes (BBHs) for alternative configurations of space- and ground-based detectors on the detection of additional GW polarizations. We employ the ppE framework for model-independent testing in GR by numerically calculating time-domain GW signals that contain all polarizations. Space-based detectors LISA, Taiji, and TianQin are used to observe massive black hole binaries (MBHBs), while ground-based detectors LIGO, Virgo, KAGRA, and the Einstein Telescope (ET) observe stellar-mass binary black holes (SBBHs), providing constraint results under different networks. We also consider the combination of multiband observations. The long-term observation of SBBHs by space-based detectors helps break the response degeneracy of ground-based detectors, thereby improving constraints. Using the Fisher information matrix (FIM), we present the constraint results of several typical mass BBHs as a function of redshift. Additionally, we discuss the potential effects of multimessenger observations on enhancing parameter constraints. Through systematic research, we comprehensively analyze the results of additional GW polarizations constrained by space- and ground-based detectors from multiple perspectives.
@@
This paper is organized as follows. In Sec. II, we introduce time-domain GW signals and ppE parameters for measuring additional polarizations within the ppE framework. In Sec. III, we review the performance parameters and alternative configurations of space- and ground-based detectors, and analyze the response functions corresponding to different modes. In Sec. IV, we explain the typical BBH sources selected in our paper and the method for calculating the signal-to-noise ratio (SNR) and FIM. In Sec. V, we present the constraint results on ppE parameters, including observations from different networks, multiband, and multimessenger observations. Finally, we summarize the results of our research in Sec. VI.
@@ II Gravitational wave signal
Figure 1: Comparison of time-domain GW waveforms under GR and ppE frameworks. The GW waveform of an MBHB with a total mass of $10^{5}\ \mathrm{M_{\odot}}$ is observed by LISA over the year before coalescence. To clearly visualize differences, we set $\alpha_{D}=0.001$, $\alpha_{Q}=96/5$, and $\alpha_{V}=\alpha_{B}=\alpha_{L}=0.15$, ensuring $\mathcal{A}_{V}=\mathcal{A}_{B}=\mathcal{A}_{L}$. (a) displays the specific time-domain GW signal. (b) and (c) show the GW waveforms from 415 days and 30 minutes before coalescence, reflecting the early and late inspiral phases, respectively. (d) illustrates the ratio of the amplitude of additional modes to the tensor modes.
In generalized modified gravity theories, metric perturbations can have up to six independent degrees of freedom, resulting in six different polarization modes: two tensor modes ($+$ and $\times$), two vector modes ($X$ and $Y$) and two scalar modes ($B$ and $L$[47]. For detectors, the observed GW strain can be described as a linear combination of different GW polarizations, expressed as
@@
The extended ppE framework is used to construct a model-independent test for GR, including all GW polarization modes [37]. The amplitude and phase of GWs can be obtained separately from the measurement of perturbations and energy evolution, and there exist simultaneous quadrupole and dipole radiation. The process of CBC includes three phases: inspiral, merger, and ringdown. Currently, the observed GW events are transient and do not include the early inspiral phase [19]. Moreover, GR has passed all current experimental tests, indicating that only tensor modes generated by quadrupole radiation are present in the non-inspiral phase. The contribution of dipole radiation in the early inspiral phase can be greater than during the merger phase [36, 48]. Therefore, it is reasonable to assume that tensor modes dominate in quadrupole radiation, while vector and scalar modes dominate in dipole radiation. Following Refs. [49, 36, 37, 47], the GW waveform under the ppE framework can be written as
@@
We only consider the dominant modes in different radiations for the amplitude, and the contribution of different radiations to the orbital angular frequency varies. Considering the influence of possible non-dominant modes, the overall evolution of the orbital angular frequency can be described by the contributions of dipole radiation and quadrupole radiation [49, 36]:
@@
To describe the model-independent GW waveform, we select five dimensionless ppE parameters to constrain parameters in alternative gravity theories. In the case of GR, $\alpha_{D,V,B,L}=0$ and $\alpha_{Q}=96/5$. To ensure our analysis remains independent of specific models, we treat these ppE parameters as independent. Some theories exhibit correlations between these parameters, potentially enhancing parameter constraints [36]. Thus, our results are conservative, and the actual constraints may exceed those presented here.
From Eq. (4), solving the evolution function $\omega(t)$ of the orbital angular frequency over time analytically is challenging. We determine $t(\omega)$ through integration::
@@
which is shown in Refs. [49, 36]. Computational methods, such as the bisection method, iteratively solve for the orbital angular frequency corresponding to a given time point. By solving point by point, we obtain $\omega(t)$. Using the above method, we input the calculated $\omega(t)$ into Eqs. (1)-(3) to derive the final GW signal.
In order to better demonstrate the impact of ppE parameters on GW waveforms significantly, we set large parameter values in Fig. 1. As shown in Fig. 1(b), there is a notable amplitude difference between the ppE waveform and the GR waveform. This discrepancy arises because the GW frequency from dipole radiation equals the orbital frequency of the BBH, whereas quadrupole radiation operates at twice that frequency. This effectively superimposes a waveform with half the frequency onto the original GR waveform, resulting in amplitude modulation at twice the wavelength. Comparing Figs. 1(b) and (c), the disparity between the ppE waveform and the GR waveform is significant when the merger is distant but diminishes as the merger approaches, consistent with our assumption in Eq. (2) to align with current GR testing outcomes. In Fig. 1(d), the ratio of additional modes to tensor modes is visually depicted, showing a gradual decrease over time, causing the ppE waveform to progressively approach the GR waveform. From Eq. (3), it can be seen that the amplitude of the tensor modes varies with the power of $2/3$ of the angular frequency, while the additional modes varies with the power of $1/3$. The rate of amplitude increase for tensor modes is significantly greater than for additional modes, resulting in the ratio $\mathcal{A}_{V}/\mathcal{A}_{T}\propto\omega^{-1/3}$ declining as angular frequency increases.
@@ III.1 Space-based detectors
Figure 2: Configuration and location of space- and ground-based detectors (not to scale). (a) and (b) depict alternative orbital configurations of LISA, Taiji-p, Taiji-c, Taiji-m, TianQin I, and TianQin II [44]. The solid arrow on the constellation plane represents the normal direction of the detector constellation plane. (c) shows the geographical locations and azimuth angles of LIGO, Virgo, KAGRA, and ET [50, 51].
@@
In the 2030s, LISA, Taiji, and TianQin are scheduled to launch triangular constellations consisting of three spacecraft. LISA and Taiji utilize heliocentric orbits, whereas TianQin employs geocentric orbits. Various alternative orbital configurations are available, as detailed in Figs. 2 and 3. LISA and Taiji, including three configurations, have a leading or trailing angle of $20^{\circ}$ relative to Earth, with a $60^{\circ}$ inclination between the constellation plane and the ecliptic plane [31, 32]. Furthermore, the normal direction of the TianQin constellation plane remains fixed, and TianQin follows a “three months on+three months off” observation scheme [33].
Variations in the constellation plane determine how GW sources appear in the detector coordinate system, thereby influencing the response to different polarizations. Additionally, detector sensitivity is directly influenced by the arm length, with LISA, Taiji, and TianQin having arm lengths of $2.5\times 10^{6}$ km, $3\times 10^{6}$ km, and $\sqrt{3}\times 10^{5}$ km, respectively. Concerning the noise in space-based detectors, we account for acceleration noise, displacement noise, and foreground noise. Acceleration noise and displacement noise contribute to the primary power spectral density (PSD), as detailed in Refs. [53, 54, 55]. Foreground noise originates from galactic binaries within the Milky Way galaxy, creating a peak in the frequency band of approximately 0.3-3 mHz, as demonstrated in our previous work [44].
@@ III.2 Ground-based detectors
In contrast to the planned space-based detectors, the ground-based detectors are currently operational, primarily comprising the LVK collaboration, which includes LIGO, Virgo and KAGRA. These are all L-shaped second-generation detectors with arm lengths of 4 km (LIGO), 3 km (Virgo), and 3 km (KAGRA), respectively [28, 29, 30]. In this paper, we consider LIGO to include three detectors: LIGO Hanford (H1) and LIGO Livingston (L1), which are operational, and the planned LIGO India (I1). Furthermore, we consider the third-generation detector ET, which features a triangular shape composed of three 10 km arms [56].
As the Earth rotates, ground-based detectors scan different sky regions, as illustrated in Fig. 3. Due to the obliquity of the ecliptic, the sky regions scanned by space- and ground-based detectors is not parallel. Ground-based detectors, unlike their space-based counterparts, are affected by various types of noise, including quantum noise, seismic noise, gravity-gradient noise, thermal noise, and others. For our study, we utilize the design performance specifications of these detectors, and the corresponding PSD can be referenced in LIGO Document T1500293 and Refs. [57, 58].

### v2

@@ Abstract
The theory of general relativity (GR) predicts the existence of gravitational waves (GWs) with two tensor modes, while alternative theories propose up to six polarization modes. In this study, we investigate constraints on GW polarization using a model-independent parametrized post-Einsteinian framework and consider both space- and ground-based detectors. By evaluating the capabilities and network performance of LISA, Taiji, TianQin, LIGO, Virgo, KAGRA, and the Einstein Telescope (ET), we analyze their respective contributions. Among space-based detectors, Taiji provides the most stringent constraints compared with LISA and TianQin. Regarding ground-based detectors, LIGO excels in vector modes while ET offers comprehensive constraints across all polarization modes. In network scenarios, LISA+TJm performs best, and ET surpasses second-generation detector combinations. Furthermore, multiband observations effectively mitigate scalar mode degeneracies thereby significantly enhancing the performance of ground-based detectors. Ultimately, combined space- and ground-based observations provide robust constraints on GW polarizations that advance tests for deviations from GR. Our findings underscore the potential of future GW missions in refining our understanding of gravitational physics through precise measurements of polarizations.
@@ I Introduction
Over the past century, general relativity (GR) has undergone rigorous testing through numerous solar system experiments and astrophysical observations, yielding no definitive evidence suggesting deviations from GR [1, 2, 3]. Despite successfully passing all experimental tests, there are still indications that GR may require extension due to the challenges from dark energy and dark matter in cosmology [4, 5]. Gravitational waves (GWs) provide a new way to test GR, especially in the strong-field regime where GWs originate from compact objects [6]. One method to test GR is by detecting additional polarizations of GWs. In GR, GWs possess only two tensor modes; however, certain alternative gravity theories predict at most six types of polarizations [7]. For instance, Brans-Dicke theory predicts an extra scalar mode [2, 8], while $f(R)$ theory and Horndeski theory propose two additional scalar modes [9, 10, 11]. Einstein-Aether theory predicts five polarization modes [12, 13], and some tensor-vector-scalar theories encompass all six polarization modes [14, 15]. Therefore, detecting additional polarizations of GWs can aid in identifying deviations from GR beyond current experimental limits and unraveling the profound nature of gravity.
@@
Since the first detection of GW by LIGO in GW150914 [16], nearly a hundred GW events generated by compact binary coalescences (CBCs) have been observed [17, 18, 19]. The analysis of these GW data has consistently demonstrated that all observed events are in agreement with GR so far [20, 21, 22, 23, 24, 25, 26]. The start of the fourth observation run (O4), is expected to lead to the detection of more GW events, offering additional opportunities for testing GR. It is crucial to use multiple detectors capable of capturing signals from various directions to effectively detect the polarization of transient GWs and enhance detection capabilities. [27]. Ground-based detectors operating in the hertz frequency band, such as LIGO [28], Virgo [29], and KAGRA [30], collaborate to form a network for measuring the polarization of GWs. For continuous GWs emitted by distant CBCs, a detector with directional change capability alone would suffice for accurate measurements [27]. Space-based GW detectors operating in the millihertz frequency band, including LISA [31], Taiji [32], and TianQin [33]. They consist of three spacecraft forming a triangular formation, orbiting either the Sun or Earth. These configurations enable the detection of additional polarizations due to the change in direction [34].
@@
Theoretically, the capability of a detector to observe additional polarizations of GWs should be assessed, along with verifying the consistency of the detected GW signal with GR predictions. One approach involves utilizing a waveform model with purely phenomenological parameters to represent potential deviations from GR and applying constraints to these parameters using observed GW data [6]. Yunes and Pretorius introduced the parametrized post-Einstein (ppE) waveform [35], derived from the post-Newtonian (PN) approximation, to effectively parametrize the impact of alternative gravity theories on non-GR polarizations. Most modified gravity theories, such as Brans-Dicke theory, massive gravity, and bimetric theory, can be characterized within the ppE framework [36]. The ppE framework cannot parametrize all possible deviations from GR, but it does offer a method for testing additional polarizations [37, 38, 39].
@@
At present, the ppE framework is being employed to impose constraints on additional polarizations for testing GR. Narikawa and Tagoshi explored the potential of advanced ground-based detectors, such as LIGO, Virgo, and KAGRA, in detecting deviations beyond GR using the ppE framework [40]. Takeda et al. estimated the accuracy of polarization determination for CBCs using the Fisher information matrix (FIM) [41, 42]. Huwyler et al. employed LISA to detect massive black hole binaries (MBHBs) and test GR by representing ppE waveforms with phase corrections only [43]. The potential of Taiji and TianQin in testing GR and constraining non-GR parameters using MBHBs has been discussed in Refs. [36, 6]. Nair et al. investigated the synergistic impact of the Einstein Telescope (ET) and preDECIGO on detector networks, demonstrating improved sensitivity within a specific frequency range [44]. Xie et al. utilized GW signals from nearly ten thousand double white dwarfs observed by the TianQin and LISA networks to detect additional polarization modes [45]. In Ref. [46], Wang and Han utilized the LISA-Taiji network to investigate the ppE parameters of deviations from the GR waveform, demonstrating a significant enhancement in the detection of polarization amplitude through combined observations.
@@
Based on our previous work [47, 48, 49], we investigate the constraints from binary black holes (BBHs) for alternative configurations of space- and ground-based detectors on the detection of additional GW polarizations. The ppE framework is employed for model-independent testing in GR through the numerical calculation of time-domain GW signals that encompass all polarizations. The space-based detectors LISA, Taiji, and TianQin are employed for the observation of MBHBs, while the ground-based detectors LIGO, Virgo, KAGRA, and ET observe stellar binary black holes(SBBHs). These different networks provide constraint results under distinct observational platforms. The combination of multiband observations is also taken into consideration. The long-term observation of SBBHs by space-based detectors effectively resolves the response degeneracy in ground-based detector, enhancing their precision. The FIM is used to present the constraint results for several typical mass BBHs based on luminosity distance. The potential impacts of multimessenger observations on improving parameter constraints are also discussed. Through a systematic research approach, we comprehensively analyze the outcomes derived from additional GW polarizations that are constrained by both space- and ground-based detectors, considering multiple perspectives.
@@
This paper is organized as follows. In Sec. II, we introduce time-domain GW signals and ppE parameters for measuring additional polarizations within the ppE framework. In Sec. III, we evaluate the performance and configurations of space and ground detectors, analyzing their response functions for various modes. In Sec. IV, we explain the typical BBH sources selected in our paper and the method for calculating the signal-to-noise ratio (SNR) and FIM. In Sec. V, we present the constraint results on ppE parameters, including observations from different networks, multiband, and multimessenger observations. Finally, we summarize the results of our research in Sec. VI.
@@ II Gravitational wave signal
Figure 1: Comparison of time-domain GW waveforms under GR and ppE frameworks. The GW waveform of an MBHB with a total mass of $10^{5}\ \mathrm{M_{\odot}}$ is observed by LISA over the year before coalescence. To highlight differences, we set $\alpha_{D}=0.001$, $\alpha_{Q}=96/5$, and $\alpha_{V}=\alpha_{B}=\alpha_{L}=0.15$, ensuring $\mathcal{A}_{V}=\mathcal{A}_{B}=\mathcal{A}_{L}$. (a) displays the specific time-domain GW signal. (b) and (c) show the GW waveforms from 415 days and 30 minutes before coalescence, reflecting the early and late inspiral phases, respectively. (d) illustrates the ratio of the amplitude of additional modes to the tensor modes.
In generalized modified gravity theories, GWs can have at most six types of polarizations: two tensor modes ($+$ and $\times$), two vector modes ($X$ and $Y$) and two scalar modes ($B$ and $L$[50]. For detectors, the observed GW strain can be described as a linear combination of different GW polarizations, expressed as
@@
The extended ppE framework is utilized to construct a model-independent test for GR, encompassing all GW polarization modes [37]. The amplitude and phase of GWs can be obtained from the metric perturbation and energy evolution, respectively [36]. The process of CBC includes three phases: inspiral, merger, and ringdown. The current observations of GW events are transient and do not encompass the early inspiral phase [19]. The successful passing of all current experimental tests in GR provides evidence for the dominance of tensor modes during the merger and ringdown phases. The contribution of dipole radiation in the early inspiral phase may exceed that during the merger phase [36, 51]. The dominance of tensor modes in quadrupole radiation and the dominance of vector and scalar modes in dipole radiation can be reasonably assumed. Following Refs. [52, 36, 37, 50], the GW waveform under the ppE framework can be written as
@@
The overall evolution of orbital angular frequency can be described by considering the combined effects of dipole and quadrupole radiation [52, 36]
@@
The model-independent GW waveform is described by selecting five dimensionless ppE parameters to constrain the parameters in alternative theories of gravity. In the case of GR, $\alpha_{D,V,B,L}=0$ and $\alpha_{Q}=96/5$. The independence of our analysis is guaranteed by treating these ppE parameters as mutually independent. Theories often show correlations among these parameters, potentially enhancing the constraints on them [36]. The findings presented here are conservative, and the actual results may be more positive.
The analytical solution for the time evolution function $\omega(t)$ of the orbital angular frequency, as indicated by Eq. 4), poses a challenging task. We determine $t(\omega)$ through integration,
@@
which is shown in Refs. [52, 36]. The bisection method and other computational techniques are employed to iteratively determine the orbital angular frequency corresponding to a given time point. By solving point by point, we obtain the value of $\omega(t)$. Using the above method, we input the calculated $\omega(t)$ into Eqs. (1$\sim$(3) to derive the final GW signal.
In order to better demonstrate the impact of ppE parameters on GW waveforms significantly, we have employed the parameters with large values in Fig. 1. As shown in Fig. 1(b), a noticeable disparity in amplitude is observed between the ppE waveform and the GR waveform. The discrepancy arises because the GW frequency resulting from dipole radiation is equivalent to the orbital frequency of the BBH, while quadrupole radiation operates at the twice frequency. Consequently, this effectively overlays a waveform with a half of the frequency onto the original GR waveform, leading to amplitude modulation at the twice wavelength. The disparity between the ppE waveform and the GR waveform is significant when the merger is far away but diminishes as the merger approaches, which aligns with our assumption in Eq. (2) and current GR testing outcomes (see Figs. 1(b) and (c)). The visual representation in Fig. 1(d) shows a gradual decrease in the ratio of additional modes to tensor modes over time, causing the ppE waveform to converge towards the GR waveform. The amplitude of tensor modes varies with an angular frequency raised to the power of $2/3$, while additional modes vary with an angular frequency raised to the power of $1/3$ according to Eq. (3). The amplitude of tensor modes increases more quickly than that of the additional modes, resulting in a declining ratio $\mathcal{A}_{V}/\mathcal{A}_{T}\propto\omega^{-1/3}$ as angular frequency increases.
@@ III.1 Space-based detectors
Figure 2: Configuration and location of space- and ground-based detectors (not to scale). (a) and (b) depict alternative orbital configurations of LISA, Taiji-p, Taiji-c, Taiji-m, TianQin I, and TianQin II [53, 54, 55, 56, 47]. The solid arrow on the constellation plane represents the normal direction of the detector constellation plane. (c) shows the geographical locations and azimuth angles of LIGO, Virgo, KAGRA, and ET [57, 58].
@@
The LISA, Taiji, and TianQin missions are planned to launch triangular constellations comprising three spacecraft in the 2030s. LISA and Taiji will adopt heliocentric orbits, while TianQin will employ geocentric orbits. Various alternative orbital configurations are available, as illustrated in Figs. 2 and 3. The LISA and Taiji missions exhibit a leading or trailing angle of $20^{\circ}$ with respect to Earth, while maintaining a $60^{\circ}$ inclination between the constellation plane and the ecliptic plane [31, 32]. The TianQin constellation plane remains fixed in its normal direction and follows an observation scheme characterized by alternating periods of “three months on+three months off” [33].
The orientation of the constellation plane affects how GW sources appear in the detector coordinate system and influences their response to different polarizations. The sensitivity of the detectors is directly affected by the length of their arms. LISA, Taiji, and TianQin have arm lengths of $2.5\times 10^{6}$ km, $3\times 10^{6}$ km, and $\sqrt{3}\times 10^{5}$ km respectively. The noise in space-based detectors includes acceleration, displacement, and foreground noise. The primary power spectral density (PSD) includes contributions from acceleration and displacement noises, as described in Refs. [60, 61, 62]. The foreground noise comes from galactic binaries in our Milky Way galaxy, creating a peak frequency band of 0.3-3 mHz [47].
@@ III.2 Ground-based detectors
Unlike the planned space-based detectors, the operational ground-based detectors primarily include the LVK Collaboration comprising LIGO (4 km), Virgo (3 km), and KAGRA (3 km) [28, 29, 30]. These detectors are all second-generation with an L-shaped configuration. In this paper, we consider LIGO as comprising three operational detectors: LIGO Hanford (H1), LIGO Livingston (L1), and the planned LIGO India (I1). The third-generation detector ET is characterized by a triangular shape with three arms, each measuring 10 km [63].
The rotation of the Earth causes ground-based detectors to scan different regions of the sky, as depicted in Fig. 3. Due to the obliquity of the ecliptic, the sky regions scanned by space- and ground-based detectors are not parallel. The performance of ground-based detectors is affected by various types of noise, including quantum noise, seismic noise, gravity-gradient noise, thermal noise, and others. In our study, we utilize the design specifications for these detectors and refer to the corresponding PSD in LIGO Document T1500293 and Refs. [64, 65].